find the equation of the straight line that passes through the intersection point between the lines x + y = 3 and 2x - 3y = 16 and perpendicular to the straight line y = 4x - 3
1
Expert's answer
2020-11-02T20:42:10-0500

x + y = 3 in the form y=mx+c

will be "y=-x+3" ........(equation 1)


2x - 3y = 16 in the form y=mx+c

will be "y=\frac{2}{3}x - \frac{16}{3}"........(equation 2)

At point of intersection (equation 1) is equal to (equation 2)


"-x+3=\frac{2}{3}x - \frac{16}{3}"


"3+\frac{16}{5} = \frac{2}{3}x+x"


"\frac{25}{3} =\frac{5}{3}x"


Hence. "x = 5"

From (equation 1) "y=-x+3"

Therefore

"y = -5 +3"

"y = -2"

Hence point of intersection is (5,-2)



For perpendicular lines the product of their gradient is -1

For line "y=4x-3" the gradient is "4"

Let the gradient of the other line be M2


M2 x 4 = -1

M2 = -1/4


"\frac{y--2}{x-5} = \frac{-1}{4}"


"4y+8=-x+5"

"4y = -x-3"


The equation of the line will be,

"y= -\frac{1}{4}x -\frac{3}{4}"